6. Let A= 0.7 0.1 0.3 0.9 Find P and D such that A= PDP-1, where D is a diagonal matrix.This is a two-parts of the same question please answer both7. This is a continuation of the previous problem. Compute lim AK ) Po, where ро 0.8 0.2
Answer #1

() And bin Given Aero9 0.3 09 of A, let For eigenvalues TA-Io 0 - 0 3 04-/- (7) (-) - 0.03 Eo 3) 063 0.7) - 09 > + 2. -0030 ?(2) let X=[] be the eigenvecter Corresponding the eigenvalue i=1, then (A-II)X1= 10:02 07-1 0.3 0:9-][39] [] -0.3 he co] [52](3) (A -0.61) X220 0.9-0.6 0.1 50296 1J-[:] los *1*1[] ]-[:] 0.3 0.14 tool Y Z 6.34 + 0.34 on equating 0.14 toly zo and 0.34tWe have O 8 1 | (4) a b e d 으 then -d -b adj P= 년 a = . adjp= Then P= I adj P IPI - 3 PE -1 - 1 야 -3 문 New Apa at - 1 개 고 - 3(5) Hence D= AS [6 86] 0.6 from (1) PAP: р => PCPAP) P= POP → (PPT) A (pp)) = ppp! c) A= PPP (2) Ans (7) foom (2) A = PDP A²(6) Similarly А Po² p = ho p[ develop P-[o] AK = Pokp =p Now lum K100 AK 11 um (PDk py) kyo = Plum ok ok j pt Koo p · Plum k(7) lim Ak ha 1 il s] 1/4 34 내 Kyo 3 1/4 ht h/s 윤 314 Nw 0 0 Um Ak KA0o h// ht / sh/s AILe 0.2 0 0 +2 1 밤 24 + 90 30 h 1/4 31